OpenStudy (anonymous):

Evaluate and leave the solution in standard notation: (2 x 104)(5.1 x 102)

OpenStudy (anonymous):

\[\Large (2\times 104)(5.1\times 102)\] does X mean multiply?

Directrix (directrix):

(2 x 10^4)(5.1 x 10^2) =============== (2 x (10^4)) * (5.1 x (10^2)) = 10 200 000

OpenStudy (anonymous):

Well Thank you!!:)